OpenStudy (anonymous):

How would I solve 5y-4y=1/2(3-1)?

hero (hero):

\(5y - 4y = \dfrac{1}{2}(3 - 1)\)

OpenStudy (anonymous):

So do I do 3-1 first?

hero (hero):

Yes, parentheses first always

OpenStudy (anonymous):

okay so then we have 5y-4y=1/2(2)

OpenStudy (anonymous):

so two times 1/2?

OpenStudy (anonymous):

5y-4y=4?

hero (hero):

Wait a minute. Are you trying to put the previous problem in point-slope form?

hero (hero):

If so, then you have entered numbers incorrectly.

OpenStudy (anonymous):

yes...oh

OpenStudy (anonymous):

Yeah I'm really bad at this

OpenStudy (anonymous):

I plugged in what you gave me on the last question and that's what I cam up with

hero (hero):

No, that's not what I said to do. You mis-interpreted me.

OpenStudy (anonymous):

oh okay so this was what the questions said: Determine the equation of the line, in point-slope form, that will get your spacecraft from Point A to Point B. Point A: (1,4) Point B: (3,5)

hero (hero):

The point-slope form is: \(y-y_1 = m(x-x_1)\) The slope is \(m = \dfrac{1}{2}\) the first point is \((x_1, y_1) = (1,4)\) Thus the equation is: \(y - 4 = \dfrac{1}{2}(x - 1)\)

OpenStudy (anonymous):

okay so aren't 5 and 3 the x variables?

hero (hero):

No, you can't call numbers variables.

OpenStudy (anonymous):

oh okay...So where does the 5 and 3 go?

hero (hero):

We don't need the 5 or 3 because they are not part of the point-slope formula

OpenStudy (anonymous):

Oh got it...So how do I finish solving?

hero (hero):

For the point slope formula, we only need the first point and the slope. We used the 5 and the 3 to calculate the slope, so it's not like we didn't use them.

OpenStudy (anonymous):

Okay...So first I need to solve y−4=1/2 (x−1) ?

hero (hero):

That's the equation of the line in point-slope form. You don't need to do anything else. That is what they were asking for.

OpenStudy (anonymous):

Okay...I think I made it harder than what it needed to be... Thank you so much ^-^

OpenStudy (anonymous):

Could you help me with the second part?

OpenStudy (anonymous):

Determine the equation of the line, in slope-intercept form, that will get your spacecraft from Point B to Point C. Point B: (3,5) Point C: (5,3)

hero (hero):

Okay, did you find the slope yet?